OpenStudy (anonymous):

write 21/8 as a mixed number. I got 2 5/8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

you are right

OpenStudy (venomblast):

Trust yourself

OpenStudy (austinl):

How many times can you put 8 into 21 without getting a fraction? 2 times right? So we now have 21-16. Therefore, \(\dfrac{21}{8}=2\dfrac{5}{8}\)

OpenStudy (austinl):

So, simply... correctomundo.

OpenStudy (anonymous):

thanks guys